# CarbonNet: How Computer Vision Plays a Role in Climate Change? Application: Learning Geomechanics from Subsurface Geometry of CCS to Mitigate Global Warming

### 09 Mar 2024 v1

@@ 1 Introduction
CCS is an important emissions reduction technology to alleviate global warming and climate change[22]. In CCS, the supercritical $CO_{2}$ is injected into the deep brine aquifers or abandoned oil reservoirs for permanent storage. However, the possible risk is the surface upheaval induced by excessive fluid pressure, which is possible to damage the structures on the ground and influence people’s lives [15]. Understanding and predicting the spatial distribution and temporal change of surface deformation is important to avoid the possible hazards. However, its realization is challenged by the large spatial scale and long time scale of the project, which makes the direct numerical simulation computationally prohibitive. As the traditional numerical simulation usually requires extensive computing resources[29, 1, 14, 9, 23], the neural networks may provide a less accurate, but more efficient and sufficient solution beneficial for the model-based decision making and planning [25]. We formulate the entire project as the workflow described in Fig. 1. The project is to predict the variation of surface vertical displacement (upheaval) with time induced by injected fluid under different subsurface geometry settings. The input are 2D images of subsurface geometry consisting of a shale layer embedded with different dip angles in two permeable layers. The output is the displacement image. The shale layer and permeable layer have large contrast in their mechanical property as stiffness $\mathbb{C}^{e}$ and flow property as permeability $\boldsymbol{\kappa}$ [11].
@@
From the ResNetUNet in static mechanics model, we continue to the LSTM and transformer model to predict the temporal change of surface displacement. The input to our algorithm is a geometry image along with five frames of displacement curves, which is in format of a short video. We then use LSTM and transformer to output predicted displacement curve (Fig. 1 transient prediction output) at the next time step. LSTM is a typical algorithm for time series predictions, and we are motivated to test transformer because of its self-attention mechanism. It provides the possibility to predict the long term dependency by contracting any two locations to a constant, so that it overcomes the issue of information loss in sequence computing in LSTM [32].
@@ 2 Related Work
Deep learning approaches have been proved to be effective on predicting the static and time-dependent dynamics of a system under the governing physics principles as partial differential equations (PDE). The current existing deep learning approaches are of two types as physics-based neural networks (PINN) [31, 17, 33] and the simulation data driven approaches[25, 26, 30, 18]. The difference is whether the direct simulation data is needed to train the model.
PINN method obtains the approximate solution by minimizing the residual in physics governing equations. In contrast, common numerical discretization methods use finite difference method (FDM) and finite element method (FEM). PINN is a mesh-free method to employ automatic differentiation to handle differential operators. The existing PINN model for coupled flow and geomechanics problem[4] is limited to one dimensional problem, and it ignores the influence of material heterogeneity. Besides, it only predicts the excess pore pressure in flow side, but not the surface displacement from geomechanics side. The other type of deep learning approaches is simulation data driven method, which takes the solution from simulation directly as training data. The end-to-end image based approaches take both input and output, from figures with the model geometry as input to the contour plot showing the distribution of different physical quantities (fluid pressure, displacement) as output. The ResNet CNN was applied to predict stress field of a 2D cantilever beam under different boundary conditions [18]. Nevertheless, it is limited to static solid mechanics problem ignoring the coupling between fluid flow and mechanics response. Besides, it only deals with homogeneous model, which is not the case in reality. The R-U-Net model with LSTM is used to predict the temporal and spatial distribution of surface displacement [25], fluid pressure and saturation in the 3D model. However, it is restricted to predict one specific geological formation setting.
In order to predict the temporal change of the surface displacement under different geological formation settings, a more advanced model is needed to capture the dynamics changing with time. Transformer and self-attention models [28], which were originally developed in NLP, shows its ability to outperform other methods by learning longer-term dependencies without recurrent connections. Transformer has been used to predict the representative dynamical systems such as 2D fluid dynamics and 3D reaction-diffusion dynamics [18]. The model shows its effectiveness to capture the varying behavior of different physical phenomenon evolution. Therefore, transformer is chosen to predict the transient behavior of surface displacement under different subsurface geometry setting in our project as well.
@@ 3 Dataset
In the section, we demonstrate the dataset used in the project. The dataset includes two parts as the static model and the transient model. The input dataset are 2D geometries of geological layers (Fig. 1 input image), which consists of two heterogeneous materials as the hard rock with high permeability and soft shale with low permeability; and the label is surface displacement contour (Fig. 1 static prediction output image) under specific loading. The geometries are sampled according to the real underground stratum. The size of the domain (length and width of the rectangle), the thickness, and the depth of shale layer is fixed, while the variable is the dip angle of shale layer. The data of surface displacement field is generated from simulation results in FEniCS (an open-source computational platform using Finite element Method (FEM)) [2]. The model setting used in the problem is depicted in Fig. 2. Data augmentation is not necessary in the project because we could generate as many training samples as we want from the physics simulation.
Figure 2: The layout of the computational problem. The end of the computational program is to capture the displacement given the configuration of the loading, material distribution and boundary conditions[3].
@@ 3.0.2 Static displacement as the label
We consider vertical displacement $u_{y}$ as the label, which is defined on each vertices of the finite element mesh [10], and can be considered as the displacement at points uniformly sampled over the domain. The vertical displacement ($u_{y}$) corresponds to the upheaval of the stratum. The FEniCS program provides the following API:
@@ 4.1 Static mechanics model
We are motivated to start from CNN models by literature [18], and further test ResNet and UNet from a similar deep learning problem to handle heterogeneity in input images [8]. The CNN and ResNet network reduces the dimension of images layer by layer. It loses information while collecting important features from input. Meanwhile, the deep neural network constructed by the ResNet can be utilized to decode the physical Eq. 6, which is a complicated procedure. Additionally, the heterogeneous distribution of material impacts the final result considerably. Inspired by the segmentation in computer vision [19], we use UNet to capture the heterogeneous distribution of the pixel/material.
@@ 4.1.2 Model: ResNetUNet
ResNetUNet is the model designed by ourselves to solve the current problem. Compared to the conventional UNet [19], we modify the encoder part by using the ResNet for downsampling the image to better capture the physical equations. The upsampling is the same as the traditional UNet [34]. The architecture of the ResNetUNet is illustrated in the Fig. 3. The special part here is the residual block, which replace the ordinary CNN layer in UNet with two consecutive residual blocks, as is shown in Fig. 4. To successfully fulfill the functionality, we need to capture both the heterogeneous features and the physics equations laws. The UNet part is to do a segmentation-like work to distinguish the shale and rock, while the ResNet part is to learn the complicated physics laws.
@@ 4.2 Transient mechanics model
In this part, we use both the LSTM and transformer to predict the evolution of vertical displacement, the interest in industry, given the specified material distribution as an image. We adopt two models, LSTM and transformer, to perform the task, which is similar to an image caption work [24, 16], except that we use the displacement as the sequence rather than the words embedding.
@@ 4.2.1 LSTM model
The combination of CNN and LSTM have been proved effective for the image caption work [24]. The CNN can be used to extract the features of the input image to feed into the LSTM network flow together with the sequence to be processed. In this project, we use a pretrained model in the static case to process the input geometry to get the features, which can embody the heterogeneous distribution of the material. The pretrained model used here is the ResNetUNet model trained in the static case, which has been proved effective for the feature extraction. Then, we add a linear layer for enhancing expressivity of the model, and processing the features further to enter the LSTM network. The linear layer has the architecture of in_features=1250, and out_feature=40.
For LSTM model, we use five steps of data points as the input to predict the next single step. Since the dimension of the sequence is 40, the number of points sampled on the ground surface, thus, we use the input size of the sequence as 40. The LSTMCell model, which includes input modulation gate, input gate, forget gate and output gate [6], in Pytorch is adopted, and it is initialized by setting the input size as 40, and the hidden size as 4. We stack four LSTMCells to construct the model. The main architecture is depicted in Fig. 5. The loss function used in this model is MSE, which is the same as the Eq. 4. Also, we use both MSE and MAE as the metrics to evaluate the model.
The architecture of this model is inspired by the image caption model of LSTM [21]. The difference lies in that we use the continuous value of the displacement as the data, rather than the softmax classification of the word embedding. Thus, the loss function for the two tasks is also distinct. Based on the accuracy requirements from the industry, the dimensions of the sequence data in our engineering application are approximately the same as the dimensions of words in natural language processing. The similarity in the two models provides the confidence to build the machine learning model based on the image caption model in NLP.
@@ 4.2.2 Transformer model
The input linear layer here has input features with size 1250, and out features with size 40, which reduces the dimensions of the output of pretrained model to 40. For the encoder part, we design another linear layer to convert the dimension from 40 to 16 to optimize the memory. The output then enters the positional encoding layer to take the order of the sequence into account. Then, we employ the nn.TransformerEncoderLayer in the Pytorch to assemble the encoder. The structure of the encoder is standardized from the attention paper [27]. The parameters we set for the encoder layer are d_model=16, nhead= 8. We stack four nn.TransformerEncoderLayer to form the encoder layer. For decoder, we use nn.TransformerDecoderLayer, which have the structure of d_model=16, nhead= 8. And we stack four of them to set up the decoder layer. At the end of the transformer layer, we add a linear mapping to map the dimension inner the decoder to the sequence dimension. The main architecture of the transformer is illustrated in Fig. 6.
@@ 5.1 Static mechanics model
We selected CNN as a baseline model because it collects important information from input images as features by filtering through all pixels. We have found the different subsurface geometry has strong correlation to the displacement images, and that motivates us to try CNN first to examine its performance. We generated a total of 10,000 samples and then selected 95% for training and 5% for validation. The mini batch size is 32 because this fits our CPU/GPU memory, and we prioritize the model convergence speed for less number of epochs. Although a smaller mini batch size could be less efficient in gradient calculation terms [5], a large batch may degrade the quality of the model because it tends to converge to sharp minima [12]. We tuned the hyperparameters including learning rate, optimizer, beta1, beta2, and weight decay to expect decrease of validation loss values versus the number of epochs. We also compare the training loss versus validation loss to avoid underfitting or overfitting problems. Therefore, the learning is tuned to be 5e-6, and the weight decay is 1e-5. We finally use Adam as our optimizer, because it requires less memory with large number of parameters. More importantly, the hyperparameters have intuitive interpretation, and it requires less tuning with low computational costs [13].
@@ 5.2 Transient mechanics model
LSTM model has shown promising potential in time series prediction in various field [20, 7]. However, the current data prediction mainly refers to the low-dimensional data, which is much lower than the case we deal with here, with dimension of 40. For the training of the LSTM model, we tune parameters including the number of LSTM cells, batch__\__size, learning__\__rate. Similarly, we use an Adam optimizer for the gradient descent. To avoid the overfitting and underfitting, we finally stack 4 LTSM cells, and we choose the batch__\__size as 1024 to facilitate the execution in GPU. The learning__\__rate is chosen as 0.001.
@@ 7 Supplementary Material
We tracked our source code using Github (Click Here), and please let us know if you could not access to it because we set it to be private (this is important because we tend to make this project to be a publication). We may need your username in order to grant access.
For video prediction, refer to the demo (Click here) to view the evolution of the displacement. We also developed a web application for our users, and the link could be found on the page of Github repo. This is an alpha version of the demo of our tool.
8 Contributions
Wei Chen wrote program of finite element analysis to generate data, designed and implemented ResNetUNet and transformer. Yunan Li designed and implemented CNN, ResNet and LSTM, conducted the performance evaluation analysis among different models. Yuan Tian conducted the literature review and provided the background knowledge. Wei Chen, Yunan Li and Yuan Tian wrote the paper and finished the poster all together.

### 12 Mar 2024 v2

@@ 1 Introduction
CCS is an important emissions reduction technology to alleviate global warming and climate changestephens2006growing . In CCS, the supercritical $CO_{2}$ is injected into the deep brine aquifers or abandoned oil reservoirs for permanent storage. However, the possible risk is the surface upheaval induced by excessive fluid pressure, which is possible to damage the structures on the ground and influence people’s lives li2013application . Understanding and predicting the spatial distribution and temporal change of surface deformation is important to avoid the possible hazards. However, its realization is challenged by the large spatial scale and long time scale of the project, which makes the direct numerical simulation computationally prohibitive. As the traditional numerical simulation usually requires extensive computing resourcesVilarrasa2010 ; Shi2013 ; Li2016 ; Fuchs2019 ; Talebian2013 , the neural networks may provide a less accurate, but more efficient and sufficient solution beneficial for the model-based decision making and planning Tang2021 . We formulate the entire project as the workflow described in Fig. 1. The project is to predict the variation of surface vertical displacement (upheaval) with time induced by injected fluid under different subsurface geometry settings. The input are 2D images of subsurface geometry consisting of a shale layer embedded with different dip angles in two permeable layers. The output is the displacement image. The shale layer and permeable layer have large contrast in their mechanical property as stiffness $\mathbb{C}^{e}$ and flow property as permeability $\boldsymbol{\kappa}$ josh2012laboratory .
@@
From the ResNetUNet in static mechanics model, we continue to the LSTM and transformer model to predict the temporal change of surface displacement. The input to our algorithm is a geometry image along with five frames of displacement curves, which is in format of a short video. We then use LSTM and transformer to output predicted displacement curve (Fig. 1 transient prediction output) at the next time step. LSTM is a typical algorithm for time series predictions, and we are motivated to test transformer because of its self-attention mechanism. It provides the possibility to predict the long term dependency by contracting any two locations to a constant, so that it overcomes the issue of information loss in sequence computing in LSTM zeyer2019comparison .
@@ 2 Related Work
Deep learning approaches have been proved to be effective on predicting the static and time-dependent dynamics of a system under the governing physics principles as partial differential equations (PDE). The current existing deep learning approaches are of two types as physics-based neural networks (PINN) Xu2020 ; Meng2020 ; Zheng2020 and the simulation data driven approachesTang2021 ; Tang2020 ; Wen2021 ; nie2020stress . The difference is whether the direct simulation data is needed to train the model.
PINN method obtains the approximate solution by minimizing the residual in physics governing equations. In contrast, common numerical discretization methods use finite difference method (FDM) and finite element method (FEM). PINN is a mesh-free method to employ automatic differentiation to handle differential operators. The existing PINN model for coupled flow and geomechanics problembekele2020deep is limited to one dimensional problem, and it ignores the influence of material heterogeneity. Besides, it only predicts the excess pore pressure in flow side, but not the surface displacement from geomechanics side. The other type of deep learning approaches is simulation data driven method, which takes the solution from simulation directly as training data. The end-to-end image based approaches take both input and output, from figures with the model geometry as input to the contour plot showing the distribution of different physical quantities (fluid pressure, displacement) as output. The ResNet CNN was applied to predict stress field of a 2D cantilever beam under different boundary conditions nie2020stress . Nevertheless, it is limited to static solid mechanics problem ignoring the coupling between fluid flow and mechanics response. Besides, it only deals with homogeneous model, which is not the case in reality. The R-U-Net model with LSTM is used to predict the temporal and spatial distribution of surface displacement Tang2021 , fluid pressure and saturation in the 3D model. However, it is restricted to predict one specific geological formation setting.
In order to predict the temporal change of the surface displacement under different geological formation settings, a more advanced model is needed to capture the dynamics changing with time. Transformer and self-attention models NIPS2017_3f5ee243 , which were originally developed in NLP, shows its ability to outperform other methods by learning longer-term dependencies without recurrent connections. Transformer has been used to predict the representative dynamical systems such as 2D fluid dynamics and 3D reaction-diffusion dynamics nie2020stress . The model shows its effectiveness to capture the varying behavior of different physical phenomenon evolution. Therefore, transformer is chosen to predict the transient behavior of surface displacement under different subsurface geometry setting in our project as well.
@@ 3 Dataset
In the section, we demonstrate the dataset used in the project. The dataset includes two parts as the static model and the transient model. The input dataset are 2D geometries of geological layers (Fig. 1 input image), which consists of two heterogeneous materials as the hard rock with high permeability and soft shale with low permeability; and the label is surface displacement contour (Fig. 1 static prediction output image) under specific loading. The geometries are sampled according to the real underground stratum. The size of the domain (length and width of the rectangle), the thickness, and the depth of shale layer is fixed, while the variable is the dip angle of shale layer. The data of surface displacement field is generated from simulation results in FEniCS (an open-source computational platform using Finite element Method (FEM)) alnaes2015fenics . The model setting used in the problem is depicted in Fig. 2. Data augmentation is not necessary in the project because we could generate as many training samples as we want from the physics simulation.
Figure 2: The layout of the computational problem. The end of the computational program is to capture the displacement given the configuration of the loading, material distribution and boundary conditionsbathe2007finite .
@@ 3.0.2 Static displacement as the label
We consider vertical displacement $u_{y}$ as the label, which is defined on each vertices of the finite element mesh ho1988finite , and can be considered as the displacement at points uniformly sampled over the domain. The vertical displacement ($u_{y}$) corresponds to the upheaval of the stratum. The FEniCS program provides the following API:
@@ 4.1 Static mechanics model
We are motivated to start from CNN models by literature nie2020stress , and further test ResNet and UNet from a similar deep learning problem to handle heterogeneity in input images charng2020deep . The CNN and ResNet network reduces the dimension of images layer by layer. It loses information while collecting important features from input. Meanwhile, the deep neural network constructed by the ResNet can be utilized to decode the physical Eq. 6, which is a complicated procedure. Additionally, the heterogeneous distribution of material impacts the final result considerably. Inspired by the segmentation in computer vision ronneberger2015u , we use UNet to capture the heterogeneous distribution of the pixel/material.
@@ 4.1.2 Model: ResNetUNet
ResNetUNet is the model designed by ourselves to solve the current problem. Compared to the conventional UNet ronneberger2015u , we modify the encoder part by using the ResNet for downsampling the image to better capture the physical equations. The upsampling is the same as the traditional UNet zhou2019d . The architecture of the ResNetUNet is illustrated in the Fig. 3. The special part here is the residual block, which replace the ordinary CNN layer in UNet with two consecutive residual blocks, as is shown in Fig. 4. To successfully fulfill the functionality, we need to capture both the heterogeneous features and the physics equations laws. The UNet part is to do a segmentation-like work to distinguish the shale and rock, while the ResNet part is to learn the complicated physics laws.
@@ 4.2 Transient mechanics model
In this part, we use both the LSTM and transformer to predict the evolution of vertical displacement, the interest in industry, given the specified material distribution as an image. We adopt two models, LSTM and transformer, to perform the task, which is similar to an image caption work tan2019phrase ; liu2021cptr , except that we use the displacement as the sequence rather than the words embedding.
@@ 4.2.1 LSTM model
The combination of CNN and LSTM have been proved effective for the image caption work tan2019phrase . The CNN can be used to extract the features of the input image to feed into the LSTM network flow together with the sequence to be processed. In this project, we use a pretrained model in the static case to process the input geometry to get the features, which can embody the heterogeneous distribution of the material. The pretrained model used here is the ResNetUNet model trained in the static case, which has been proved effective for the feature extraction. Then, we add a linear layer for enhancing expressivity of the model, and processing the features further to enter the LSTM network. The linear layer has the architecture of in_features=1250, and out_feature=40.
For LSTM model, we use five steps of data points as the input to predict the next single step. Since the dimension of the sequence is 40, the number of points sampled on the ground surface, thus, we use the input size of the sequence as 40. The LSTMCell model, which includes input modulation gate, input gate, forget gate and output gate braun2018lstm , in Pytorch is adopted, and it is initialized by setting the input size as 40, and the hidden size as 4. We stack four LSTMCells to construct the model. The main architecture is depicted in Fig. 5. The loss function used in this model is MSE, which is the same as the Eq. 4. Also, we use both MSE and MAE as the metrics to evaluate the model.
The architecture of this model is inspired by the image caption model of LSTM soh2016learning . The difference lies in that we use the continuous value of the displacement as the data, rather than the softmax classification of the word embedding. Thus, the loss function for the two tasks is also distinct. Based on the accuracy requirements from the industry, the dimensions of the sequence data in our engineering application are approximately the same as the dimensions of words in natural language processing. The similarity in the two models provides the confidence to build the machine learning model based on the image caption model in NLP.
@@ 4.2.2 Transformer model
The input linear layer here has input features with size 1250, and out features with size 40, which reduces the dimensions of the output of pretrained model to 40. For the encoder part, we design another linear layer to convert the dimension from 40 to 16 to optimize the memory. The output then enters the positional encoding layer to take the order of the sequence into account. Then, we employ the nn.TransformerEncoderLayer in the Pytorch to assemble the encoder. The structure of the encoder is standardized from the attention paper vaswani2017attention . The parameters we set for the encoder layer are d_model=16, nhead= 8. We stack four nn.TransformerEncoderLayer to form the encoder layer. For decoder, we use nn.TransformerDecoderLayer, which have the structure of d_model=16, nhead= 8. And we stack four of them to set up the decoder layer. At the end of the transformer layer, we add a linear mapping to map the dimension inner the decoder to the sequence dimension. The main architecture of the transformer is illustrated in Fig. 6.
@@ 5.1 Static mechanics model
We selected CNN as a baseline model because it collects important information from input images as features by filtering through all pixels. We have found the different subsurface geometry has strong correlation to the displacement images, and that motivates us to try CNN first to examine its performance. We generated a total of 10,000 samples and then selected 95% for training and 5% for validation. The mini batch size is 32 because this fits our CPU/GPU memory, and we prioritize the model convergence speed for less number of epochs. Although a smaller mini batch size could be less efficient in gradient calculation terms bengio2012practical , a large batch may degrade the quality of the model because it tends to converge to sharp minima keskar2016large . We tuned the hyperparameters including learning rate, optimizer, beta1, beta2, and weight decay to expect decrease of validation loss values versus the number of epochs. We also compare the training loss versus validation loss to avoid underfitting or overfitting problems. Therefore, the learning is tuned to be 5e-6, and the weight decay is 1e-5. We finally use Adam as our optimizer, because it requires less memory with large number of parameters. More importantly, the hyperparameters have intuitive interpretation, and it requires less tuning with low computational costs kingma2014adam .
@@ 5.2 Transient mechanics model
LSTM model has shown promising potential in time series prediction in various field sagheer2019time ; cao2019financial . However, the current data prediction mainly refers to the low-dimensional data, which is much lower than the case we deal with here, with dimension of 40. For the training of the LSTM model, we tune parameters including the number of LSTM cells, batch__\__size, learning__\__rate. Similarly, we use an Adam optimizer for the gradient descent. To avoid the overfitting and underfitting, we finally stack 4 LTSM cells, and we choose the batch__\__size as 1024 to facilitate the execution in GPU. The learning__\__rate is chosen as 0.001.
@@ 7 Supplementary Material
Get access to our source code in Github (Click Here).
For time series prediction, refer to the demo (Click here) to view the evolution of the displacement.